1. Lovely-Tech Company produces printers. In January, the company produced 3600 printers per week. The company has 25 direct labor employees. Each employee works 40 hours a week with an hourly wage of $20.00. The material cost of the firm per week is $15,000.00. The overhead expense incurred in operating the firm is $5,000 per week.

(1) Determine the weekly multi-factor productivity of the Lovely-Tech Company in January.

(Hint: Total input = Labor cost + Material cost + Overhead)

(Labor cost =number of employees * working hours a week* wage)

(2) In February, Lovely-Tech Company produced 3350 printers per week and the material cost per week is $18,000. If the labor cost and overhead remain the same, what is growth rate (in percentage) in weekly multifactor productivity from January to February?
Answer #1

Multi Factor Productivity = Output/Input

Here, Output = Number of Printers Produced

Input = Total cost = Labor Cost + Material Cost + Overhead Cost

For January -

Input = 3600

Output = 25*40*20 + 15000 + 5000 = 40000

Multi Factor Productivity = 3600/40000 = 0.09

For February -

Input = 3350

Output = 25*40*20 + 18000 + 5000 = 43000

Multi Factor Productivity = 3350/43000 = 0.077907

Growth in Productivity = (0.077907 - 0.09)/0.09

Growth in Productivity = -0.13437 = -13.44% (Productivity has been decreased in February)